# TA-KAND: Two-stage Attention Triple Enhancement and U-KAN based Diffusion For Few-shot Knowledge Graph Completion

### v1

@@ Abstract
Knowledge Graphs (KGs), thanks to their concise and efficient triple-based structure, have been widely applied in intelligent question answering, recommender systems and other domains. However, the heterogeneous and multifaceted nature of real-world data inevitably renders the distribution of relations long-tailed, making it crucial to complete missing facts with limited samples. Previous studies mainly based on metric matching or meta learning, yet they either fail to fully exploit neighborhood information in graph or overlook the distributional characteristics of contrastive signals. In this paper, we re-examine the problem from a perspective of generative representation and propose a few-shot knowledge graph completion framework that integrates two-stage attention triple enhancer with U-KAN based diffusion model. Extensive experiments on two public datasets show that our method achieve new state-of-the-art results.
Note: This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
@@ 1 Introduction
Knowledge Graphs like Freebase [1] and Wikidata [2] have been widely introduced into diverse applications [3, 4]. A typical Knowledge Graph (KG) $\mathcal{G}$ consists of a large collection of factual triples $\mathcal{T}=\{(h,r,t)\in\mathcal{E}\times\mathcal{R}\times\mathcal{E}\}$, where $\mathcal{E}$ and $\mathcal{R}$ denote the entity set and relation set. Despite their massive scale, KGs remain inherently incomplete due to the diversity of source and the incessant update of dynamic message. A common solution is Knowledge Graph Embedding (KGE), which projects all compositions into low-dimensional vectors to discover unobserved links. However, the long-tailed distribution of relations often limits the effectiveness of this method. Therefore, Few-shot Knowledge Graph Completion (FKGC), which infers missing facts with limited information, has emerged as a critical research topic.
@@
Recently, generative models like Diffusion [12] enable us to estimate the distribution of real-world data, which inspire a conceptualization of the FKGC problem from a generative representation perspective. We view the triple embedding enhancement as a conditional encoding procedure, and regard learning the distribution over positive and negative sample pairs as the denoising objective that the diffusion model incrementally converges to. Based on this, an intuitive conjecture is that conditional encoding capable of extracting positive and negative features will exhibit superior alignment with the synthesized output, which facilitates the capture of valid triple semantic rules.
Motivated by this assumption, in this paper, we propose a framework TA-KAND that integrates Two-stage Attention Triple Enhancement and Diffusion Process. Triple enhancement is employed to produce better conditional information, and diffusion process is utilized to simulate distribution. To improve the accuracy and non-linear fitting capacity of denoise module, we investigate the incorporation of Kolmogorov-Arnold Network (KAN) units into U-Net to establish an updated architecture. Experiments on NELL and Wiki demonstrate our introduced framework outperforms existing methods. Additionally, we devise several model variants to empirically corroborate our assumption.
Fig. 1: Illustration of TA-KAND model architecture.
@@ 2 Background
Task formulation. For a KG $\mathcal{G}$, we select some relations from $\mathcal{R}$ to form three disjoint task sets $\mathcal{R}_{train}$, $\mathcal{R}_{valid}$ and $\mathcal{R}_{test}$, the remaining subset constructs the background KG $\mathcal{G}^{\prime}$. Given a relation $r\in\mathcal{R}_{train}$ with its own $K$-shot support set $\mathcal{S}_{r}=\{(h_{i},t_{i})\}_{i=1}^{K}$ and query set with $N$ elements, $\mathcal{Q}_{r}=\{(h_{j},t_{j},\mathcal{C}_{h_{j},r})\}_{j=1}^{N}$. The goal of FKGC is to train a model which can rank the true tail $t_{j}$ higher than other candidates in $\mathcal{C}_{h_{j},r}$. $\mathcal{R}_{valid}$ and $\mathcal{R}_{test}$ are defined in the same way as $\mathcal{R}_{train}$, they are used to validate and test the learned model.
@@ 3 Methodology
In this section, we introduce TA-KAND, which consists of three components: (1) Two-Stage Attention Triple Enhancer improves entity-pair embeddings while distilling the semantic knowledge of task-specific relations, (2) Diffusion Learner uses U-KAN as its backbone to model the distribution of positive and negative latent triple rules, (3) The Matcher leverages the generated latent compression representation to infer the correct tail in the query. The overall architecture is illustrated in Fig. 1.
3.1 Two-Stage Attention Triple Enhancer
In a KG, the intrinsic characteristics of both entities and relations can be implicitly or explicitly reflected through neighborhood context [7, 14]. When given a triple of few-shot task for relation $r$, we take its head entity $h$ as a target, and denote its local neighborhood as $\mathcal{N}_{h}=\{(r_{i},e_{i})|(h,r_{i},e_{i})\in\mathcal{G^{\prime}}\}$. For each neighbor, its overall representation $\boldsymbol{n_{i}}\in\mathbb{R}^{d}$ should be learned,
$\boldsymbol{n}_{i}=\sigma(W_{n}[\boldsymbol{r_{i}}||\boldsymbol{e_{i}}]+b_{n})$ (1)
where $\boldsymbol{r_{i}},\boldsymbol{e_{i}}\in\mathbb{R}^{d}$ are the embeddings of $r_{i},e_{i}$ pre-trained on $\mathcal{G}^{\prime}$ with TransE, $d$ is the embedding dimension, $||$ means concatenation, $W_{n}$ and $b_{n}$ are shared learnable parameters, $\sigma$ denotes ReLU. Next, we treat the task relation $r$ in a triple as a function of its head and tail, calculate its relevance to each neighbor relation $r_{i}$, and employ an attention mechanism to produce the aggregated neighborhood representation $\boldsymbol{h}_{nbr}$.
$\boldsymbol{r}=\boldsymbol{h}^{\mathsf{T}}W\boldsymbol{t}+b$ (2)
$\boldsymbol{h}_{nbr}=Attention(Q=W_{r}\boldsymbol{r},K=\boldsymbol{r_{i}},V=\boldsymbol{n_{i}})$ (3)
@@ 3.1 Two-Stage Attention Triple Enhancer
where $\boldsymbol{h},\boldsymbol{t}\in\mathbb{R}^{d}$ are head and tail embedding in TransE, $W,W_{r}\in\mathbb{R}^{d\times d}$ and $b\in\mathbb{R}^{d}$ are trainable parameters. Afterward, We couple the original embedding $\boldsymbol{h}$ and $\boldsymbol{h}_{nbr}$ to generate the first-stage attention enhanced head representation $\boldsymbol{h_{r}}$:
$\boldsymbol{h_{r}}=\sigma(W_{1}\boldsymbol{h}+W_{2}\boldsymbol{h}_{nbr})$ (4)
where $W_{1},W_{2}\in\mathbb{R}^{d\times d}$ are learnable parameters. The above operations also hold for tail entity $t$.
@@
Beyond relation roles, the semantics of entities also be considered, so we acquire the refined neighborhood information $\boldsymbol{h}_{nbr}^{\prime}$ by calculating the similarity between $\boldsymbol{t_{r}}$ and each neighbor tail $\boldsymbol{e_{i}}$. This procedure again employs the attention mechanism, followed by the same coupling operation used in the first stage.
$\boldsymbol{h}_{nbr}^{\prime}=Attention(Q=W_{e}\boldsymbol{t_{r}},K=\boldsymbol{e_{i}},V=\boldsymbol{n_{i}})$ (5)
$\boldsymbol{\tilde{h}}=\sigma(W_{1}\boldsymbol{h}+W_{2}\boldsymbol{h}_{nbr}^{\prime})$ (6)
@@
Here, $W_{e}\in\mathbb{R}^{d\times d}$ is a trainable matrix, and the second-stage attention process is applied to tail $t$ analogously. As a result, we obtain the two-stage attention enhanced entities $\boldsymbol{\tilde{h}}$ and $\boldsymbol{\tilde{t}}$.
Although previous approaches have emphasized learning high-quality pairs [6, 7], they neglect the explicit extraction of task relation. Therefore, we adopt an aggregate-disentangle-compress paradigm to derive the representation of task relation. Specifically, we first aggregate the enhanced head and tail entities within the $K$-shot support set into paired representations, which are disentangled directly and fed to Bi-LSTM to generate task relation descriptions. The compression operation yields a relation prototype by weighted summation over the $K$ descriptors.
$\begin{gathered}\boldsymbol{p_{j}}=[\boldsymbol{p_{j}^{1}||p_{j}^{2}}]=W_{4}\sigma(W_{3}[\tilde{\boldsymbol{h_{j}}}||\tilde{\boldsymbol{t_{j}}}]+b_{p})\\ \boldsymbol{p_{j}^{\prime}}=\text{Bi-LSTM}(\boldsymbol{p_{j}})\ \ \ (j=1,2,...,K)\\ \boldsymbol{\tilde{r}}=W_{scale}(\sum_{j}\omega_{j}\boldsymbol{p_{j}^{\prime}})\ \ \ \ \omega_{j}={\exp{\langle\boldsymbol{p_{j}^{\prime}},\boldsymbol{c}\rangle}\over\sum_{j}\exp{\langle\boldsymbol{p_{j}^{\prime}},\boldsymbol{c}\rangle}}\end{gathered}$ (7)
where $W_{3},W_{4},W_{scale}\in\mathbb{R}^{d\times d}$ are learnable, $\boldsymbol{c}$ means the cell state output in Bi-LSTM, and $\langle\cdot,\cdot\rangle$ is the dot product.
3.2 Diffusion Learner
@@
Diffusion Process is used to learn the latent triple rule representation. In forward process, Gaussian noise is progressively injected into $\boldsymbol{z_{0}}$, which consists of the enhanced entity pairs from both the support set $\mathcal{C}_{r}=\{(\boldsymbol{\tilde{h_{i}}},\boldsymbol{\tilde{t_{i}}})\}_{i=1}^{K}$ and the negative set $\mathcal{C}_{r}^{-}=\{(\boldsymbol{\tilde{h_{i}}},\boldsymbol{\tilde{t_{i}}}^{-})\}_{i=1}^{K}$ constructed by corrupting the tail entities. In denoise process, we integrate Kolmogorov–Arnold Networks (KANs) into the bottleneck of the U-Net backbone, resulting in a novel architecture for noise prediction. Concretely, the residual-downsampling blocks are employed for feature extraction and dimensional compression. At the bottleneck, the compressed features are first flatten into a token sequence via a learnable linear projection, then forward into a KAN layer followed by layer normalization. Finally, residual-upsampling blocks are performed to restore the feature dimensionality to its original scale. Conditional injection is implemented through two strategies: The task relation representation, head-tail enhanced embeddings in supprt and negative set with their corresponding labels (label is True when a triple is in support set, otherwise is False) are fused into the downsampling/upsampling network with FiLM technique [15]. In the inserted KAN module, only the encoded timestep is supplied.
@@ 3.4 Model Training
Given a specific task relation, we randomly sample some pairs to construct its positive query set $\mathcal{Q}_{r}^{+}$. For each pair in $\mathcal{Q}_{r}^{+}$, we pollute the tail and construct negative query set $\mathcal{Q}_{r}^{-}=\{(h_{q},t_{q}^{-})|(h_{q},t_{q})\in\mathcal{Q}_{r}^{+}\}$, where $t_{q}^{-}\in\{\mathcal{C}_{h_{q},r}\setminus t_{q}\}$. Our optimization object is to ensure the score of a positive query in $\mathcal{Q}_{r}^{+}$ higher than a negative query in $\mathcal{Q}_{r}^{-}$. Additionally, in Diffusion Learner, the noise added to the initial embedding $\boldsymbol{z_{0}}$ must be predicted accurately. Overall, the loss function $\mathcal{L}$ is defined as follows,
$\mathcal{L}=\sum_{r}\{\sum_{\mathcal{Q}_{r}^{+},\mathcal{Q}_{r}^{-}}\max[0,\gamma+\mathcal{M}(h_{q},r,t_{q}^{-})\\ -\mathcal{M}(h_{q},r,t_{q})]+\mathbb{E}_{t,z_{t}}||\epsilon_{\theta}-\epsilon_{t}||_{2}^{2}\}$ (9)
@@ 4.1 Experimental Settings
Datasets. We conduct experiments on two standard benchmark datasets, NELL and Wiki [5]. In both datasets, relations with less than 500 but more than 50 triples are selected as few-shot relations. Following the original setting [5], train/valid/test set are 51/5/11 in NELL and 133/16/34 in Wiki. More details are shown in Table 1.
Comparision Methods. To evaluate our method, we choose seven FKGC methods as baselines (Table 2). We omit further comparison with traditional KGE models, as they have been empirically outperformed by some FKGC approches [7].
Inplementation Details. We perform 5-shot KG completion task for all methods.
Dataset #Ent #Rel #Triple #Tasks
NELL 68,545 358 181,109 67
Wiki 4,838,244 822 5,829,240 183
Table 1: Statistics of datasets. Each column represents the number of entities, relations, triples and task relations.
Methods NELL Wiki
MRR Hits@10 Hits@5 Hits@1 MRR Hits@10 Hits@5 Hits@1
GMatching [5] $0.176$ $0.294$ $0.233$ $0.113$ $0.263$ $0.387$ $0.337$ $0.197$
FSRL [6] $0.153$ $0.319$ $0.212$ $0.073$ $0.158$ $0.287$ $0.206$ $0.097$
MetaR [10] $0.261$ $0.437$ $0.350$ $0.168$ $0.221$ $0.302$ $0.264$ $0.178$
FAAN [7] $0.279$ $0.428$ $0.364$ $0.200$ $0.341$ $0.463$ $0.395$ $0.281$
GANA [11] $0.344$ $0.517$ $0.437$ $0.246$ $0.351$ $0.446$ $0.407$ $0.299$
NP-FKGC [8] $0.460$ $0.494$ $0.471$ $0.437$ $0.503$ 0.668 0.599 $0.423$
ReCDAP [9] 0.505 0.528 0.506 0.493 0.510 $0.577$ $0.539$ 0.474
TA-KAND(Ours) 0.778 0.807 0.797 0.759 0.685 0.735 0.714 0.657
Table 2: Results of 5-shot FKGC on NELL and Wiki.
@@
A1 0.205 0.259 0.222 0.176
A2 0.452 0.458 0.452 0.441
B1 0.697 0.724 0.713 0.681
B2 0.730 0.761 0.749 0.714
TA-KAND 0.778 0.807 0.797 0.759
Table 3: Results of model variants on NELL dataset. Bold numbers denote the best performance.
The triple embeddings are initialized with the pre-trained TransE model and further fine-tuned during training. Their dimensionality is set 100 and 50 for NELL and Wiki, respectively. We fix the margin $\gamma$ to 1.0 and use the Adam optimizer [16] with an initial learning rate of $1e^{-3}$ for NELL and $1e^{-4}$ for Wiki to update parameters. The Diffusion Learner uses 1000 timesteps and the scheduler adopts the setting employed in GLIDE [17]. For efficient computation, the two-stage attention enhancement is realized using Flash Attention [18] with a single head and a scaling factor of 1.0. All experiments were conducted on a single NVIDIA A800.

### v2

@@ Abstract
Knowledge Graphs have become fundamental infrastructure for applications such as intelligent question answering and recommender systems due to their expressive representation. Nevertheless, real-world knowledge is heterogeneous, leading to a pronounced long-tailed distribution over relations. Previous studies mainly based on metric matching or meta learning. However, they often overlook the distributional characteristics of positive and negative triple samples. In this paper, we propose a few-shot knowledge graph completion framework that integrates two-stage attention triple enhancer with U-KAN based diffusion model. Extensive experiments on two public datasets show significant advantages of our methods.
Note: The method proposed in this paper, along with a related framework of a similar nature, is currently under submission. Therefore, the presentation of figures and experimental data is limited at this stage. The complete manuscript and full experimental results will be formally released once the review decisions are finalized!
@@ 1 Introduction
Knowledge Graphs like Freebase [1] and Wikidata [2] have been widely introduced into diverse applications [3, 4]. A typical Knowledge Graph (KG) $\mathcal{G}$ consists of a large collection of factual triples $\mathcal{T}=\{(h,r,t)\in\mathcal{E}\times\mathcal{R}\times\mathcal{E}\}$, where $\mathcal{E}$ and $\mathcal{R}$ denote the entity set and relation set. Despite their massive scale, KGs remain inherently incomplete due to the diversity of source and the incessant update of dynamic message. While Knowledge Graph Embedding (KGE) projects all compositions into low-dimensional vectors to discover unobserved links, the long-tailed distribution of relations limits its effectiveness. Therefore, Few-shot Knowledge Graph Completion (FKGC), which infers missing facts with limited information, has emerged as a critical research topic.
@@
In this paper, we propose a framework TA-KAND that integrates Two-stage Attention Triple Enhancement and Diffusion Process. Triple enhancement is employed to produce better conditional information, and diffusion process is utilized to simulate distribution. To improve the accuracy and non-linear fitting capacity of denoise module, we investigate the incorporation of Kolmogorov-Arnold Network (KAN) units into U-Net to establish an updated architecture. Experiments on NELL and Wiki demonstrate our introduced framework outperforms existing methods.
@@ 2 Background
Task formulation. For a KG $\mathcal{G}$, we select some relations from $\mathcal{R}$ to form three disjoint task sets $\mathcal{R}_{train}$, $\mathcal{R}_{valid}$ and $\mathcal{R}_{test}$, the remaining subset constructs the background KG $\mathcal{G}^{\prime}$. To imitate few-shot link prediction, we sample a relation $r\in\mathcal{R}_{train}$ with its own $K$-shot support set $\mathcal{S}_{r}=\{(h_{i},t_{i})\}_{i=1}^{K}$ and query set with $N$ elements, $\mathcal{Q}_{r}=\{(h_{j},t_{j}/\mathcal{C}_{h_{j},r})\}_{j=1}^{N}$. Here, $t_{j}$ is the ground-truth tail entity and $\mathcal{C}_{h_{j},r}$ denotes the candidate set, where each candidate is an entity in $\mathcal{E}$ selected according to type constraints. The goal of FKGC is to train a model which can rank the true tail $t_{j}$ higher than other candidates in $\mathcal{C}_{h_{j},r}$. $\mathcal{R}_{valid}$ and $\mathcal{R}_{test}$ are defined in the same way as $\mathcal{R}_{train}$, they are used to validate and test the learned model after sufficient training.
@@ 3 Methodology
In this section, we introduce TA-KAND, which consists of three components: (1) Two-Stage Attention Enhancer improves entity-pair embeddings while distilling the semantic knowledge of task-specific relations, (2) Diffusion Processor uses U-KAN as its backbone to model the distribution of positive and negative latent triple rules, (3) The Scorer leverages the generated latent compression representation to infer the correct tail in the query.
3.1 Two-Stage Attention Enhancer
Within a knowledge graph, both entities and relations reveal their underlying semantics through interactions with neighboring nodes and edges[7, 14]. When given a triple of few-shot task for relation $r$, we take its head entity $h$ as a target, and denote its local neighborhood as $\mathcal{N}_{h}=\{(r_{i},n_{i})|(h,r_{i},n_{i})\in\mathcal{G^{\prime}}\}$. For each neighbor, we first learn its representation. Next, we treat the task relation $r$ in a triple as a bilinear function of its head and tail, calculate its relevance to each neighbor relation $r_{i}$, and employ an attention mechanism to produce the aggregated neighborhood representation $\boldsymbol{h}_{nbr}$.
$\boldsymbol{e}^{nbr}_{i}=f(\boldsymbol{r_{i}},\boldsymbol{n_{i}})$ (1)
$\boldsymbol{r}=\text{Bilinear}(\boldsymbol{h},\boldsymbol{t})$ (2)
$Q_{1}=W_{Q}^{1}\boldsymbol{r},\ \ K_{1}=W_{K}^{1}\boldsymbol{r_{i}},\ \ V_{1}=W_{V}^{1}\boldsymbol{e}^{nbr}_{i}$ (3)
$\boldsymbol{h}_{nbr}=softmax(Q_{1}K_{1}^{T})V_{1}$ (4)
@@ 3.1 Two-Stage Attention Enhancer
where $\boldsymbol{h},\boldsymbol{t}\in\mathbb{R}^{d}$ are head and tail embedding in TransE, $W_{Q}^{1},W_{K}^{1},W_{V}^{1}\in\mathbb{R}^{d\times d}$ and $b\in\mathbb{R}^{d}$ are trainable parameters. Afterward, We couple the original embedding $\boldsymbol{h}$ and $\boldsymbol{h}_{nbr}$ to generate the first-stage attention enhanced head representation $\boldsymbol{h_{r}}$. The above operations also hold for tail entity.
@@
Beyond relation roles, the semantics of entities also be considered, so we acquire the refined neighborhood information $\boldsymbol{h}_{nbr}^{\prime}$ by calculating the similarity between $\boldsymbol{t_{r}}$ and each neighbor tail $\boldsymbol{n_{i}}$. This procedure again employs the attention mechanism, followed by the same coupling operation used in the first stage.
$Q_{2}=W_{Q}^{2}\boldsymbol{t}_{r},\ \ K_{2}=W_{K}^{2}\boldsymbol{n}_{i},\ \ V_{2}=W_{V}^{2}\boldsymbol{e}^{nbr}_{i}$ (5)
$\boldsymbol{h}_{nbr}^{\prime}=softmax(Q_{2}K_{2}^{T})V_{2}$ (6)
@@
Here, $W_{Q}^{2},W_{K}^{2},W_{V}^{2}\in\mathbb{R}^{d\times d}$ is a trainable matrix, and the second-stage attention process is applied to tail $t$ analogously. As a result, we obtain the two-stage attention enhanced entities $\boldsymbol{\tilde{h}}$ and $\boldsymbol{\tilde{t}}$. Moreover, previous approaches neglect the explicit extraction of task relation, so we adopt an aggregate-disentangle-compress paradigm to derive the representation of task relation.
3.2 Diffusion Processor
@@
Diffusion Process is used to learn the latent triple rule representation. In forward process, Gaussian noise is progressively injected into $\boldsymbol{z_{0}}$, which consists of the enhanced entity pairs from both the support set $\mathcal{C}_{r}=\{(\boldsymbol{\tilde{h_{i}}},\boldsymbol{\tilde{t_{i}}})\}_{i=1}^{K}$ and the negative set $\mathcal{C}_{r}^{-}=\{(\boldsymbol{\tilde{h_{i}}},\boldsymbol{\tilde{t_{i}}}^{-})\}_{i=1}^{K}$ constructed by corrupting the tail entities. In denoise process, we integrate Kolmogorov–Arnold Networks (KANs) into the bottleneck of the U-Net backbone, resulting in a novel architecture for noise prediction. Concretely, the residual-downsampling blocks are employed for feature extraction and dimensional compression. At the bottleneck, the compressed features are first flatten into a token sequence via a learnable linear projection, then forward into a KAN layer followed by layer normalization. Finally, residual-upsampling blocks are performed to restore the feature dimensionality to its original scale. Conditional injection is implemented through two strategies: The task relation representation, head-tail enhanced embeddings in support and negative set with their corresponding labels (label is True when a triple is in support set, otherwise is False) are fused into the downsampling/upsampling network with FiLM technique [15]. In the inserted KAN module, only the encoded timestep is supplied.
@@ 3.4 Model Training
Given a specific task relation, we randomly sample some pairs to construct its positive query set $\mathcal{Q}_{r}$. For each pair in $\mathcal{Q}_{r}$, we pollute the tail and construct negative query set $\mathcal{Q}_{r}^{-}=\{(h_{q},t_{q}^{-})|(h_{q},t_{q})\in\mathcal{Q}_{r}^{+}\}$. Our optimization object is to ensure the score of a positive query in $\mathcal{Q}_{r}$ higher than a negative query in $\mathcal{Q}_{r}^{-}$. Additionally, in Diffusion Processor, the noise added to the initial embedding $\boldsymbol{z_{0}}$ must be predicted accurately. Overall, the loss function $\mathcal{L}$ is defined as follows,
$\mathcal{L}=\sum_{r}\{[\gamma-score(Q_{r}^{-})+score(Q_{r})]_{+}+\mathbb{E}_{t,z_{t}}||\epsilon_{\theta}-\epsilon_{t}||_{2}^{2}\}$ (8)
Here, $score(Q_{r})$ and $score(Q_{r}^{-})$ denote the scores of positive and negative samples in the query set, respectively. $_{+}$ is hinge loss.
@@ 4.1 Experimental Settings
Datasets. We conduct experiments on two standard benchmark datasets, NELL and Wiki [5]. In both datasets, relations with less than 500 but more than 50 triples are selected as few-shot relations. Following the original setting [5], train/valid/test set are 51/5/11 in NELL and 133/16/34 in Wiki.
Comparison Methods. To evaluate our method, we choose seven FKGC methods as baselines (Table 1). We omit further comparison with traditional KGE models, as they have been empirically outperformed by some FKGC approaches [7].
Implementation Details. We perform 5-shot KG completion task for all methods. The triple embeddings are initialized with the pre-trained TransE model and further fine-tuned during training. Their dimensionality is set 100 and 50 for NELL and Wiki, respectively. We fix the margin $\gamma$ to 1.0 and use the Adam optimizer [16] with an initial learning rate of $1e^{-3}$ for NELL and $1e^{-4}$ for Wiki to update parameters. The Diffusion Learner uses 1000 timesteps and the scheduler adopts the setting employed in GLIDE [17]. For efficient computation, the two-stage attention enhancement is realized using Flash Attention [18] with a single head and a scaling factor of 1.0. All experiments were conducted on a single NVIDIA A800.
Methods NELL
MRR Hits@10 Hits@5 Hits@1
GMatching [5] $0.176$ $0.294$ $0.233$ $0.113$
FSRL [6] $0.153$ $0.319$ $0.212$ $0.073$
MetaR [10] $0.261$ $0.437$ $0.350$ $0.168$
FAAN [7] $0.279$ $0.428$ $0.364$ $0.200$
GANA [11] $0.344$ $0.517$ $0.437$ $0.246$
NP-FKGC [8] $0.460$ $0.494$ $0.471$ $0.437$
ReCDAP [9] 0.505 0.528 0.506 0.493
TA-KAND(Ours) >0.7 >0.7 >0.7 >0.7
Table 1: Results of 5-shot FKGC on NELL. The specific data is currently pending further testing and tuning. Due to manuscript review restrictions, we are temporarily withholding the data on our Wiki. We will present the complete tables once the results are officially published. This section merely serves to indicate that our model has achieved state-of-the-art (SOTA) performance.
@@
V1 0.205 0.259 0.222 0.176
V2 0.697 0.724 0.713 0.681
V3 0.730 0.761 0.749 0.714
Table 2: Results of model variants on NELL dataset.